# Learning interactions between Rydberg atoms

### v1

@@ I Introduction
One of the frontiers of quantum characterization and verification is Hamiltonian learning, i.e. inferring the experimentally realized Hamiltonian through measurements. In realizable cases, the Hamiltonian can be directly reconstructed from knowledge of the steady states of the system [21, 22, 23, 24], the wavefunction following a quantum quench  [25, 26] or the wavefunction following the application of single-qubit pulses  [27]. However, the realization of such methods is not always feasible in an arbitrary quantum system due to their reliance on knowledge of the amplitudes of the quantum state.
@@
Given the ongoing success of machine learning (ML) in quantum physics [37, 38], especially in quantum state tomography [39, 40, 41, 42, 43, 44] and phase classification [45, 46, 47, 48, 49, 50], it is promising to address the outstanding challenges of Hamiltonian learning with neural networks. This direction has already been fruitful [24, 51, 52, 53, 54, 55, 56, 57] and outperformed Bayesian methods in a direct comparison [55]. While these neural-network-based methods are able to quickly predict many Hamiltonian parameters at once with low error based on a user-defined form of measurements, they require accurate numerical simulations of the quantum system under consideration and, therefore, do not scale up to large two-dimensional (2D) systems.
@@
In Sec. II, we introduce the GNN architecture and its training procedure, and describe in more detail the ultracold Rydberg platform, which inspired the learning task. There, we also establish a bijective correspondence between the correlation functions and the interaction parameters of the Hamiltonian, which provides a theoretical foundation for our learning algorithm. In Sec. III, we demonstrate the scalability of the GNN and the dependence of its prediction error on data representation. We discuss the critical system size that needs to be included in the training set to achieve scalability and show how the GNN prediction error depends on the number of snapshots used to estimate correlators. We furthermore compare the GNN performance to that of a multi-layer perceptron (MLP). In Sec. IV, we conclude by discussing open problems and the next directions unlocked by our Hamiltonian learning approach.
@@ II Methods
To emulate the TFIM on a Rydberg-atom platform, we employ the Density Matrix Renormalization Group (DMRG) algorithm to solve the TFIM, introduced in Sec. II.1. In Sec. II.2, we prove the bijective relation between the spin exchange couplings and the spin-spin correlation functions. This bijective relation ensures, for instance, that NN correlation functions suffice to uniquely determine NN couplings defining the Hamiltonian. Then, in Sec. II.3, we introduce the key ML techniques used to predict the atomic displacements from the DMRG local magnetization and spin correlators, as well as the metrics used to judge the predictive power of the neural network.
II.1 Physical Hamiltonian
The Hamiltonian under consideration is the 2D transverse-field Ising model (TFIM)
@@ II.1 Physical Hamiltonian
where $i<j$ runs over all atomic positions while avoiding double-counting of pairs, $C_{6}/\hbar\simeq 5.42\times 10^{6}\,\text{rad}\cdot\mu\text{s}^{-1}\cdot\mu\text{m}^{6}$ is the dipole-dipole interaction term originating from the van der Waals interaction for Rubidium with principal quantum number $n=70$ [63], with $\hbar$ the reduced Planck constant. The parameter $\Omega$ stands for the transverse Ising field (Rabi frequency), $\delta$ for the detuning field, whilst $\hat{\sigma}^{z}$ and $\hat{\sigma}^{x}$ denote the corresponding Pauli matrices. The distance $R_{ij}$ describes the pairwise distance between two atoms $i$ and $j$ sitting in the array. Hence, the interaction term (first in Eq. (1)) is an ‘all-to-all’ interaction, and it is responsible for the Rydberg blockade whenever the Rabi frequency $\hbar\Omega\ll C_{6}/R^{6}_{ij}\equiv J_{i,j}$. The two-level system is encoded in the Rydberg states, typically $s$ orbitals. In the following, we will set $\hbar=1$.

### v2

@@ I Introduction
One of the frontiers of quantum characterization and verification is Hamiltonian learning, i.e. inferring the experimentally realized Hamiltonian through measurements. In realizable cases, the Hamiltonian can be directly reconstructed from knowledge of the steady states of the system [21, 22, 23, 24], the wavefunction following a quantum quench  [25, 26] or following the application of single-qubit pulses  [27]. However, the realization of such methods is not always feasible in an arbitrary quantum system due to their reliance on knowledge of the amplitudes of the quantum state.
@@
Given the ongoing success of machine learning (ML) in quantum physics [37, 38], especially in quantum state tomography [39, 40, 41, 42, 43, 44] and phase classification [45, 46, 47, 48, 49, 50, 51], it is promising to address the outstanding challenges of Hamiltonian learning with neural networks. This direction has already been fruitful [24, 52, 53, 54, 55, 56, 57, 58] and outperformed Bayesian methods in a direct comparison [56]. While these neural-network-based methods are able to quickly predict many Hamiltonian parameters at once with low error based on a user-defined form of measurements, they require accurate numerical simulations of the quantum system under consideration and, therefore, do not scale up to large two-dimensional (2D) systems.
@@
In Sec. II, we describe in more detail the ultracold Rydberg platform, which inspired the learning task. There, we also introduce a bijective correspondence between the correlation functions and the interaction parameters of the Hamiltonian, which provides a theoretical foundation for our learning algorithm. This proof is laid out in Appendix A. Finally, we describe the learning algorithm itself and introduce the GNN architecture and its training procedure. In Sec. III, we demonstrate the scalability of the GNN and the dependence of its prediction error on data representation. We also show how the GNN prediction error depends on the number of snapshots used to estimate correlators. We furthermore compare the GNN performance to that of a multi-layer perceptron (MLP). In Sec. IV, we conclude by discussing open problems and the next directions unlocked by our Hamiltonian learning approach.
@@ II Methods
We start this section by describing briefly in Sec. II.1 the Rydberg-atom platform for quantum simulation, which inspired our learning task, as well as the Hamiltonian studied in this work. In Sec. II.2, we outline a proof, detailed in Appendix A, that demonstrates a bijective relation between spin exchange couplings and spin-spin correlation functions. This bijective relation ensures, for instance, that NN correlation functions suffice to uniquely determine NN couplings defining the Hamiltonian, and therefore the relative distances of all the NN Rydberg atoms on the lattice. Then, in Sec. II.3, we introduce the key ML techniques used to predict the atomic displacements from the DMRG local magnetization and spin correlators, as well as the metrics used to judge the predictive power of the neural network.
II.1 Rydberg platform and the studied Hamiltonian
@@
In this work, we solve the Hamiltonian learning task for the 2D transverse-field Ising model (TFIM)
@@
where $i<j$ runs over all atomic positions while avoiding double-counting of pairs, $C_{6}/\hbar\simeq 5.42\times 10^{6}\,\text{rad}\cdot\mu\text{s}^{-1}\cdot\mu\text{m}^{6}$ is the dipole-dipole interaction term originating from the van der Waals interaction for Rubidium with principal quantum number $n=70$ [64], with $\hbar$ the reduced Planck constant. The parameter $\Omega$ stands for the transverse Ising field (Rabi frequency), $\delta$ for the detuning field, whilst $\hat{\sigma}^{z}$ and $\hat{\sigma}^{x}$ denote the corresponding Pauli matrices. The distance $R_{ij}$ describes the pairwise distance between any two atoms $i$ and $j$ sitting in the array. Hence, the interaction term (first term in Eq. (1)) covers all pairs of Rydberg atoms over the whole array, and is in essence a long-range interaction which is responsible for the Rydberg blockade whenever the Rabi frequency $\hbar\Omega\ll C_{6}/R^{6}_{ij}\equiv J_{i,j}$. The two-level system is encoded in the Rydberg states, typically $s$ orbitals. In the following, we will set $\hbar=1$.
To emulate the TFIM on a Rydberg-atom platform, we employ the Density Matrix Renormalization Group (DMRG) algorithm [68, 69] to solve the TFIM. We provide more details on the DMRG calculations in Appendix E.